# Application-Driven Value Alignment in Agentic AI Systems: Survey and Perspectives

### v1

@@ 4.1.2 Dataset Construction Methodologies
Constructing high-quality value alignment datasets is critical for reliably evaluating whether different systems adhere to appropriate value frameworks. To support diverse application scenarios, researchers have primarily adopted three data construction approaches: manual construction, automatic construction, and hybrid construction that integrate human and machine-generated content.
@@
  • Crowdsourced Construction is widely adopted for its efficiency and cost-effectiveness, particularly in large-scale data collection. Many datasets—such as Social Bias Frames [294], Flames [281], and BOLD [271]—leverage crowdsourcing platforms, often with multiple rounds of quality control to ensure annotation reliability. These datasets not only scale well but also capture diverse perspectives from annotators of varied cultural and social backgrounds, enabling more comprehensive assessments of model behavior.
  • Integration and Refinement of Existing Datasets involve aggregating and modifying multiple public sources. Notable examples include IndieValueCatalog [283], DecodingTrust [278], MFQ [285], Persona Bias [279], and GlobalOpinionQA [282], all constructed through careful curation and consolidation of open-source data. This approach reduces data acquisition costs while leveraging existing high-quality resources, thereby improving research efficiency and enabling broader evaluation coverage.
Automatic Construction. With the rapid advancement of large language models (LLMs), LLM-based automatic dataset construction has become a promising approach. This method typically involves designing prompt templates to guide the model in generating question–answer pairs or scenario-based judgments, followed by manual verification to ensure data quality. Representative datasets developed using this approach include NaVAB [290], KorNAT [284], , CultureSPA [275], and DailyDilemmas [277]. Compared to manual construction, it enables the creation of more diverse and context-rich data with significantly less human effort, allowing for large-scale dataset generation in a relatively short time. However, concerns remain regarding the quality of model-generated content. LLMs may produce hallucinated outputs, and their inherent biases can result in misalignment with human values. Consequently, ensuring the reliability and value alignment of automatically constructed datasets still requires careful oversight and validation.

### v2

@@ B.1 Dataset Construction Methodologies
Constructing high-quality datasets for value alignment or evaluation is critical for reliably assessing whether systems adhere to appropriate value frameworks. To support diverse application scenarios, researchers have primarily adopted three data construction approaches: manual construction, automatic construction, and hybrid construction that integrates human and machine-generated content.
@@
Table B-II: Datasets for Value Alignment and Evaluation.
AC: Automatic Construction, CC: Crowdsourced Construction, HC: Hybrid Construction, EDC: Expert-Driven Construction, IRED: Integration and Refinement of Existing Datasets. BJ: Binary Judgments, RT: Ranking Tasks, MCQ: Multiple-Choice Questions, OEQ: Open-Ended Questions, RSQ: Rating Scale Questions, A: Alignment. E: Evaluation
Datasets Macro Level Meso Level Micro Level Construction Methodologies Question Formats Type Size
BBQ(Bias Benchmark for QA) [353] EDC MCQ E 58.4k
BeaverTails [354] IRED OEQ E 30.2k
BOLD [355] $C$ OEQ E 23.6k
CBBQ(Chinese Bias Benchmark Dataset) [356] HC MCQ E 106.5k
CDEval [357] AC MCQ E 2.9k
CDial-Bias [358] $C$ OEQ E 28k
CORGI-PM [359] IRED BJ, OEQ A, E 32.9k
CrowS-Paris $C$ BJ E 1.5k
CultureSPA [360] AC MCQ E 13k
Cvalues [361] $C$ MCQ, OEQ E 6.4k
DailyDilemmas [362] AC MCQ E 2.7k
DecodingTrust [363] IRED MCQ E 152.4k
DEFSurveySim [364] HC MCQ, RSQ E 1.1k
EEC(Equity Evaluation Corpus) [365] EDC BJ E 8.6k
ETHICS [366] $C$ MCQ, BJ, RT E 134.4k
Flames [367] $C$ OEQ A, E 1k
German Credit Data [368] IRED BJ, RT A 1k
GlobalOpinionQA [369] IRED MCQ A 2.5k
HofstedeCulturalDimensions [370] HC MCQ A, E 0.2k
IndieValueCatalog [371] IRED MCQ A 93.2k
KorNAT [372] AC MCQ A, E 10k
LLMGlobe [373] HC RSQ, OEQ E 37.6k
LaWGPT [374] IRED OEQ A 300k
MFQ(Moral Foundations Questionnaire) [375] IRED MCQ E 11k
Moral Beliefs [376] AC MCQ, BJ, OEQ E 0.5k
Moral Integrity Corpus [377] AC RSQ, OEQ E 38k
Moral Stories [378] $C$ MCQ A, E 12k
MoralExceptQA [379] EDC RT E 0.2k
NaVAB [380] AC MCQ, BJ E 63.8k
Persona Bias [365] IRED MCQ, BJ E 110.6k
PkuSafeRLHF [381] AC OEQ A 476.4k
ProgressGym [382] HC RSQ, OEQ A 1.4k
SafeSora [383] AC OEQ A 14.7k
Scruples [384] EDC MCQ, BJ E 657k
Social Bias Frames [385] $C$ RT, OEQ E 150k
Social Chemistry 101 [386] $C$ BJ, OEQ E 292k
StereoSet [387] $C$ RT, OEQ E 4.2k
ToxiGen [388] AC OEQ E 6.5k
UnQover [389] IRED BJ, OEQ E 2713k
ValueNet [390] $C$ BJ, OEQ A 21.3k
WikiGenderBias [391] IRED OEQ E 45k
WinoBias [392] EDC MCQ E 3.1k
WinoGender [393] EDC BJ, OEQ E 0.7k
@@
  • Crowdsourced Construction is widely adopted for its efficiency and cost-effectiveness, particularly in large-scale data collection. Many datasets—such as Social Bias Frames [385], Flames [367], and BOLD [355]—leverage crowdsourcing platforms, often with multiple rounds of quality control to ensure annotation reliability. These datasets not only scale well but also capture diverse perspectives from annotators of varied cultural and social backgrounds, thereby improving research efficiency and supporting a wider range of alignment and evaluation scenarios.
  • Integration and Refinement of Existing Datasets involves aggregating and modifying multiple public sources. Notable examples include IndieValueCatalog [371], DecodingTrust [363], MFQ [375], Persona Bias [365], and GlobalOpinionQA [369], all constructed through careful curation and consolidation of open-source data. This approach reduces data acquisition costs while leveraging existing high-quality resources, thereby improving research efficiency and enabling broader evaluation coverage.
Automatic Construction. With the rapid advancement of LLMs, LLM-based automatic dataset construction has become a promising approach. This method typically involves designing prompt templates to guide the model in generating question–answer pairs or scenario-based judgments, followed by manual verification to ensure data quality. Representative datasets developed using this approach include NaVAB [380], KorNAT [372], , CultureSPA [360], and DailyDilemmas [362]. Compared to manual construction, it enables the creation of more diverse and context-rich data with significantly less human effort, allowing for large-scale dataset generation in a relatively short time. However, concerns remain regarding the quality of model-generated content. LLMs may produce hallucinated outputs, and their inherent biases can result in misalignment with human values. Consequently, ensuring the reliability and value alignment of automatically constructed datasets still requires careful oversight and validation.